### v1

# Whom We Trust, What We Fear: COVID-19 Fear and the Politics of Information

Abstract.
The COVID-19 pandemic triggered not only a global health crisis but also an infodemic, an overload of information from diverse sources influencing public perception and emotional responses. In this context, fear emerged as a central emotional reaction, shaped by both media exposure and demographic factors. In this study, we analyzed the relationship between individuals’ self-reported levels of fear about COVID-19 and the information sources they rely on, across nine source categories, including medical experts, government institutions, media, and personal networks. In particular, we defined a score that ranks fear levels based on self-reported concerns about the pandemic, collected through the Delphi CTIS survey in the United States between May 2021 and June 2022. We found that both fear levels and information source usage closely follow COVID-19 infection trends, exhibit strong correlations within each group (fear levels across sources are strongly correlated, as are patterns of source usage), and vary significantly across demographic groups, particularly by age and education. Applying causal inference methods, we showed that the type of information source significantly affects individuals’ fear levels. Furthermore, we demonstrated that information source preferences can reliably match the political orientation of U.S. states. These findings highlight the importance of information ecosystem dynamics in shaping emotional and behavioral responses during large-scale crises.
Fear, Source of information, COVID-19, Causal inference, Online survey, United States
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 32st SIGKDD Conference on Knowledge Discovery and Data Mining; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Causal networksccs: Applied computing Health informatics
1. Introduction
The COVID-19 pandemic represented a global health crisis of unprecedented scale and complexity. At the same time, it generated an infodemic—a rapid and overwhelming surge of information from a wide range of sources (Solomon et al., 2020). As individuals sought clarity and guidance, they turned to news media, social platforms, government announcements, and public health organizations. The credibility, consistency, and framing of these sources played a critical role in shaping public understanding of the threat. Additionally, they influenced emotional responses across populations (Jones et al., 2021; Lu et al., 2021).
In this context, fear emerged as one of the most prominent psychological responses to the pandemic (Coelho et al., 2020). This emotion was not driven solely by the objective health threat, but also by the volume, tone, and perceived trustworthiness of the information consumed. The impact of these information sources on fear varied across different demographics. Differences in age, education, gender, and socioeconomic status significantly shaped how individuals selected, interpreted, and trusted various sources (Nino et al., 2021). These disparities highlight the necessity of accounting for both the nature of the information disseminated and the characteristics of the recipients when evaluating the psychological impact of information exposure.
Furthermore, information played a dual role during the crisis: it served both as a practical resource for guiding health-related decisions and as a symbolic indicator of political identity and civic orientation. The choice to trust or dismiss specific sources frequently mirrored broader ideological affiliations, shaping not only individual behaviors but also the level of collective trust in public institutions. Against this backdrop, the risks posed by misinformation—particularly when it stemmed from institutional or otherwise authoritative sources—became especially salient. Perceptions of inconsistency, politicization, or unreliability in official communications had the potential to undermine not only adherence to public health guidance but also confidence in democratic institutions. Understanding how various demographic groups navigated this complex informational environment is, therefore, critical for designing communication strategies that bolster both public health outcomes and civic trust in future crises.
Figure 1. Schematic representation of our study.
1.1. Contributions
Our main contributions are as follows—summarized in Figure 1. First, using statistical tests and correlation analyses on U.S. data, we quantify how different sources of information influence individuals’ levels of fear related to COVID-19. Second, we investigate how demographic factors—specifically age and education—not only affect emotional responses but also significantly shape information-seeking behavior and trust, revealing notable patterns across population groups. Third, through a causal inference model (Pearl, 2010), we compare the relative impact of age, education, and information source on fear, finding that information source is the most influential driver. Finally, building on this, we apply K-means clustering (Hartigan et al., 1979) to group U.S. states based on their residents’ reliance on different types of information. Remarkably, the resulting clusters exhibit a strong alignment with state-level political orientation—specifically, whether a state tends to lean Democratic or Republican. Collectively, our findings underscore the critical role of information sources in shaping public responses to large-scale crises. They also highlight the dangers of misinformation—particularly when disseminated by institutional or authoritative entities—which can distort public perception, exacerbate fear, and undermine trust in democratic institutions.
The remainder of this article is organized as follows. Section 2 reviews prior research on fear related to COVID-19, identifying key findings and existing gaps. Section 3 describes the data sources and the construction of the indices used in our analysis. In Section 4, we examine how fear levels vary according to the information sources individuals rely on. Section 5 introduces a causal inference model to identify the variables most strongly associated with fear. In Section 6, we analyze patterns of information reliance across the U.S. population and apply clustering techniques to group states accordingly. Finally, Section 7 summarizes our key findings and discusses their broader implications.
2. Related work
A growing body of research—predominantly based on survey data and often employing standardized tools such as the Fear of COVID-19 Scale (FCV-19S)—has explored the psychological impact of the COVID-19 pandemic, particularly the prevalence and determinants of fear. Developed by Ahorsu et al. (Ahorsu et al., 2022) in Iran, the FCV-19S has enabled researchers to quantify and compare fear levels across populations and has since been validated in various countries, including Spain (Mart´ınez-Lorca et al., 2020), Turkey (Gokseven et al., 2022), and India (Doshi et al., 2021), proving to be a reliable and consistent measure.
Several studies highlight how fear was shaped by both media exposure and interpersonal communication. For instance, Ruobing (Li, 2022) found that heightened fear was linked to increased media use and more danger-focused conversations, creating a feedback loop that sustained anxiety. However, individuals also employed strategies such as avoidance, counterarguing, and reappraisal, with counterarguing emerging as an effective method to reduce fear and encourage continued engagement. Complementing this, a study in Lebanon demonstrated that increased media exposure was positively associated with adherence to preventive behaviors, mediated by perceived knowledge and fear, emphasizing the complex role of media in shaping health-related actions during the pandemic (Melki et al., 2022).
Gender, age, health status, and socio-economic factors also played significant roles in shaping fear responses. Studies from diverse regions—including Chile (Cerda and Garc´ıa, 2022), Turkey (Gokseven et al., 2022), and India (Doshi et al., 2021)—consistently found that women, individuals with chronic illnesses, those with limited knowledge of the pandemic, and people facing economic hardship or living alone experienced greater levels of fear. Conversely, older age, higher education, and greater COVID-19 knowledge appeared to mitigate these fears. These findings underscore the importance of context-specific mental health strategies, particularly for vulnerable demographic groups. Notably, average fear scores varied significantly across population subgroups. For example, Doshi et al. (Doshi et al., 2021) found that while more than half of participants in India reported low levels of fear, certain groups—including women, married individuals, and healthcare workers—exhibited significantly higher fear levels.
Broader reviews and national surveys further enrich this picture. Quadros et al. (Quadros et al., 2021), in a scoping review of early pandemic studies, reported the FCV-19S ranging from 18.1% to 45.2%, particularly among women, healthcare workers, and those with pre-existing mental health concerns. In the U.S., Warren et al. (Warren et al., 2021) highlighted the role of anxiety sensitivity—especially regarding physical symptoms—in driving COVID-19-related fear, which in turn was linked to heightened anxiety and depression. Adding a longitudinal perspective, a 14-month study tracked fear trajectories over time, finding that while fear can be protective by promoting preventive behaviors, excessive fear contributes to anxiety and depression, illustrating the delicate balance between adaptive and maladaptive emotional responses during the pandemic (Mertens et al., 2023).
Overall, these findings emphasize that fear of COVID-19 was not uniformly experienced but varied significantly by individual characteristics and broader contextual factors. As Bhattacharjee et al. (Bhattacharjee and Acharya, 2020) noted, the pandemic’s mental health toll was widespread, with many Americans expressing fear of infection, death, and economic loss early on. This collective distress highlights the urgent need for culturally sensitive, targeted mental health interventions, particularly during public health crises.
Question Possible responses
(D2) What is your age? (1) 18-34 years, (2) 35-64 years, (3) 65+ years
(D8) What is the highest degree or level of school you have completed? (1) Lower than high school, (2) High school / Bachelor’s degree, (3) Postgraduate
(I5) In the past 7 days, from which of the following sources have you received news and information about COVID-19? (1) Doctors, (2) Scientists, (3) CDC, (4) Government, (5) Politicians, (6) Journalists, (7) Friends and family, (8) Religious leaders, (9) None of the above (other sources not listed here).
(G1) How much do you worry about catching COVID-19? (1) A great deal, (2) A moderate amount, (3) A little, (4) Not at all.
Table 1. Selected questions and responses from the Delphi US CTIS survey (Delphi Group, CMU, 2020), with labels as in the original survey.
To summarize, numerous studies have utilized the FCV-19S to examine pandemic-related fear, consistently identifying factors such as media exposure, anxiety sensitivity, and demographic variables—including age and gender—as key predictors. These investigations reveal robust associations between fear, anxiety, and depression across diverse populations, including older adults, healthcare workers, and students. Systematic reviews further highlight the high prevalence of fear during the pandemic and underscore the critical need for targeted mental health interventions.
However, most prior research has been cross-sectional and largely focused on isolated individual predictors, with limited attention to the combined effects of demographics and information sources on fear. Additionally, geographic and political influences remain underexplored. Our study addresses these gaps by employing advanced statistical and causal inference methods to evaluate how demographics and sources of information jointly shape fear responses. Moreover, by clustering U.S. states according to patterns of information source usage and linking these clusters to political orientation, we offer a comprehensive, integrated perspective on the social and contextual determinants of COVID-19 fear.
It is important to note that the FCV-19S was not employed in this study, as it is based on a specific survey (Ahorsu et al., 2022). Instead, as outlined in Section 3, our analysis is based on a distinct dataset collected through an alternative survey. Consequently, direct comparisons with results derived from the FCV-19S are not feasible.
3. Data
Throughout this work, we relied on multiple sources of COVID-19 data, including official surveillance data from the COVID-19 Data Hub (Guidotti and Ardia, 2020; Guidotti, 2022) and survey-based data from the Delphi US CTIS (Delphi Group, CMU, 2020), which contains information on demographics, the types of information sources used by individuals, and self-reported levels of fear. All datasets used in this study are open-source and publicly available. Aggregated data from the Delphi US CTIS survey can be accessed as described in the Availability of materials and data section.
We primarily used the COVID-19 Data Hub (Guidotti and Ardia, 2020; Guidotti, 2022) as our source of surveillance data. Specifically, we used daily new fatalities and confirmed cases—computed from cumulative counts—as well as the total U.S. population. These indicators were used to estimate the daily number of official infections via Sybil (Baccega et al., 2024). Specifically, this data allows us to reconstruct the dynamics of a Susceptible-Infected-Recovered-Deceased-Susceptible (SIRDS) compartmental model with reinfections over the study period. In this model, the infected compartment represents the estimate of the official number of infections—we assumed an average recovery time of 14 days to compensate for the lack of recovery data, which is largely unavailable. While the Data Hub provides additional information, such as vaccination rates and hospital statistics, these were not relevant to our study.
Since the spring of 2020, Facebook has collaborated with both Carnegie Mellon University (CMU) and the University of Maryland (UMD) through its Data for Good program (Facebook, 2020) to promote large-scale, daily surveys developed by these institutions (Astley et al., 2021; Fan et al., 2020; Kreuter et al., 2020). The Delphi Group at CMU administers the U.S. COVID-19 Trends and Impact Survey (Delphi US CTIS) (Delphi Group, CMU, 2020), which collects approximately 50,000 responses per day from Facebook users in the United States. This survey includes self-reported information on symptoms, testing, vaccination, behaviors, and other COVID-19-related topics. Access to the data was obtained through an agreement with CMU, UMD, and Facebook, as described in the Ethical Declaration section. The full set of survey questions in the U.S. is available at (Delphi Group, Carnegie Mellon University, [n. d.]).
Specifically, we computed daily aggregated information considering questions about age, education, source of information, and infection-induced fear—a summary of the selected questions and their corresponding response options is provided in Table 1. Importantly, these daily aggregates were calculated using the weights provided in the Delphi US CTIS microdata files. These weights were designed to address the differences between Facebook users and the overall US population, as well as to correct for biases associated with non-response (Salomon et al., 2021). For clarity and simplicity, throughout the remainder of the paper, we refer to responses to D2 as Age, D8 as Education, I5 as Source, and G1 as Fear. For Source, users could select multiple options, whereas for Age, Education, and Fear responses are mutually exclusive. Therefore, we accounted for all possible combinations of the nine response options in Source, totaling 256—also comprising the None of the above response, which is used in mutual exclusion with all the other possible answers.
In the remainder of the paper, we will focus on the relationships between Age, Education, and Source with respect to COVID-19–induced Fear, employing the notation presented in Table 2. Our analysis covers the period from May 2021 to June 2022 across all U.S. states, utilizing data from digital platforms such as Facebook, which combine online activity with self-reported information.
Normalization. This step primarily facilitates the later examination of correlations and comparisons among variables. By grouping respondents based on Age, Education, or not grouping at all—denoted by variable vv—, it is straightforward to derive the normalized quantities $\hat{\sigma}_{s}$:
(1) $\hat{\sigma}_{s}(v)=\frac{\sigma_{s}(v)}{\sum_{i\in\mathcal{S}}\sigma_{i}(v)},$
where we dropped the time dependence for ease of notation. Similarly, we define the normalized Fear measure $\hat{\varphi}_{s,f}$ for different levels of self-reported fear, applying an analogous procedure.
Singleton information sources. Data on information usage is provided in aggregated form, where each $s\in\mathcal{S}$ represents a possible combination of the considered information sources. However, to investigate the utilization of individual sources, we extrapolate the proportion of respondents using each singleton information source. Specifically, the average number of people using a specific singleton source $\hat{\sigma}_{s^{\prime}}$ is given by
(2) $\hat{\sigma}_{s^{\prime}}(v)=\sum_{\forall u\in\mathcal{S}:s^{\prime}\in u}\hat{\sigma}_{u}(v),$
where $s^{\prime}$ are the singletons information sources from $\mathcal{S}$. Please note that, although $\hat{\sigma}_{s^{\prime}}$ is computed starting from the normalized variable $\hat{\sigma}_{s}$, the sum of $\hat{\sigma}_{s^{\prime}}$ over all singletons $s^{\prime}$ might exceed 1 since individual values of $\hat{\sigma}_{u}$ are counted once for each $s^{\prime}\in u$.
Fear score. By assigning weights $w_{f}$ to each fear level, we define the fear score as
(3) $\hat{\varphi}_{s}(v)=\sum_{f\in\mathcal{F}}w_{f}\cdot\hat{\varphi}_{s,\ f}(v),$
where $w_{f}\in\{1,\ 0.34,\ -0.34,\ -1\}$. This formulation ensures that responses such as A great deal are counterbalanced by Not at all, with a similar compensatory relationship between A moderate amount and A little.
Symbol Description
$\mathcal{S}$ Set of 256 combinations of Source [^1] In detail, $\mathcal{S}=\big{\{}${Doctors}, {Scientists}, {CDC}, $\dots$, {None of the above}, {Doctors, Scientists}, {Doctors, CDC}, $\dots\big{\}}$ is a set containing 256 subsets.
$s^{\prime}$ Singleton information sources from $\mathcal{S}$, e.g., $\{\text{{Doctors}}\}$ and $\{\text{{Scientists}}\}$
$\mathcal{F}$ Set of self-reported fear levels $f\in\mathcal{F}$, where $\mathcal{F}=\left\{1,\ 2,\ 3,\ 4\right\}$
$\sigma_{s}$ Avg. number of respondents selecting the combination $s\in\mathcal{S}$
$\sigma_{s^{\prime}}$ Avg. number of respondents selecting singleton $s^{\prime}\in\mathcal{S}$
$\varphi_{s,\ f}$ Avg. number of respondents who self-reported a fear level $f\in\mathcal{F}$, selecting the combination $s\in\mathcal{S}$
$\varphi_{s}$ Avg. fear score among respondents selecting the combination $s\in\mathcal{S}$
$\phi_{s^{\prime}}$ Estimate of avg. fear score for singletons $s^{\prime}\in\mathcal{S}$
$v$ Demographic grouping: Age, Education, or no grouping
$t$ Time variable $t\in\left[1,\ T\right]$, where $T$ denotes the length of the time series, with one observation per day
Table 2. Notation used in the paper.
Fear score of singleton information sources. In case of Fear, we cannot directly apply the approach used in Equation 2 to obtain the average fear score from singleton source responses. This limitation arises because, for each $s\in\mathcal{S}$, the distribution of Fear across the selected sources is unknown. For example, $s=\{$CDC, Politicians, Journalists}\} denotes the average fear score across all U.S. states among respondents who selected CDC, Politicians, and Journalists as their information sources. However, this aggregated value does not reveal how much each source contributes to the self-reported fear—this could be primarily driven by Journalists, equally influenced by all three sources, or dominated by a different source within the combination.
To address this, and estimate $\phi_{s^{\prime}}$, we employ linear models to disentangle the contribution of each source to the overall average fear score as
(4) $\min_{w_{s^{\prime}}|s}\sum_{t\in T}\left(\hat{\varphi}_{s}(v,\ t)-\sum_{s^{\prime}\in s}w_{s^{\prime}|s}\cdot\hat{\varphi}_{s^{\prime}}(v,\ t)\right)^{2},$
where $w_{s^{\prime}}|s$ represents time-independent weights assigned to each singleton source $s^{\prime}$ in the combination ss—this optimization problem is solved for each $s\in S$. These weights are estimated by minimizing the squared difference between the observed average fear score of the combination $s$, denoted $\hat{\varphi}_{s}(v,\ t)$, and the weighted sum of the average fear scores of its singleton sources, $\hat{\varphi}_{s^{\prime}}(v,\ t)$. Specifically, we use $\hat{\varphi}_{s^{\prime}}(v,\ t)$—the fear score among respondents who selected $s^{\prime}$—as a proxy to estimate the weights. In this way, for each $\hat{\varphi}_{s}(v)$, we determine the contribution of each singleton $s^{\prime}\in s$. Using these weights, we then compute $\phi_{s^{\prime}}$ as
(5) $\phi_{s^{\prime}}(v)=\sum_{s\in\mathcal{S}}w_{{s^{\prime}}|s}\cdot\hat{\varphi}_{s^{\prime}}(v)$
4. Statistical analysis of fear
Figure 2. Temporal evolution of the singleton fear scores $\phi_{s^{\prime}}$—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. To reduce noise, official infections and $\phi_{s^{\prime}}$ were smoothed using a two-month rolling average.
Employing the definitions of the previous section, we investigate the drivers behind the evolution of the nationwide U.S. fear score $\phi_{s^{\prime}}$. As a starting point, we examine whether Fear differs across the various information sources. In particular, considering the nine distributions derived from the singleton information sources $s^{\prime}$, we first applied the Shapiro–Wilk test (Shapiro and Wilk, 1965) to assess the normality of each distribution, which indicated that none of the distributions follow a normal distribution. Subsequently, to evaluate whether these distributions differ significantly from one another—including both intra-group comparisons (i.e., across the nine singleton information sources within a given Age or Education group, as well as at the population level) and inter-group comparisons (i.e., comparing the nine singleton information source across different Age or Education groups)—, we performed the Kolmogorov–Smirnov test (Massey Jr, 1951)—it considers differences across the entire distribution—and the Mann–Whitney U test (Nachar et al., 2008)—which is primarily sensitive to differences in central tendency. The consistent significance across comparisons suggests meaningful variation in fear score distributions across both intra-group and inter-group levels, underscoring the influence of demographic and informational heterogeneity.
Figure 3. Spearman correlation coefficients among the time series shown in Figure 2.
In light of these findings, Figure 2 presents the temporal evolution of the singleton fear scores $\phi_{s^{\prime}}$, alongside the official number of infections across the U.S. Notably, Religious leaders and Family and friends emerge as significant outliers, consistently associated with lower fear scores [^2] Fear score spanning values around zero reflects design choice of assigning weights that counterbalance the associated responses in the survey, as defined in Equation 3.. Moreover, the fear score $\phi_{s^{\prime}}$ closely mirrors fluctuations in official infection numbers, indicating a strong correlation between the two—as shown in Figure 3, where we can notice that $\phi_{s^{\prime}}$ is more sensitive to relative increases in case numbers than to absolute levels.
Some sources show even greater sensitivity. The coefficient of variation [^3] The coefficient of variation is defined as the standard deviation divided by the absolute value of the mean, providing a measure of relative variability. is particularly high for Scientists ($\approx 3$) and the CDC ($\approx 1$), whereas all other sources remain below $0.5$, indicating moderate variability. This pattern illustrates how fear intensifies with the rising visibility of COVID-19 cases, dominating public discourse. The moderate to high variation highlights its strong dependence on fluctuations in official case numbers.
Among all sources, it is noteworthy that Family and friends and Journalists slightly anticipate phase transitions in the evolution of infections: the signal $\phi_{s^{\prime}}$ precedes both the initial rise in case numbers during the outbreak and the subsequent decline at the peak for each of the two major waves—corresponding to the Delta and Omicron variant surges (Dattani and Rodés‑Guirao, 2024). This pattern may reflect an anticipatory response based on early information or heightened awareness within these social and expert networks.
In addition, singleton fear scores exhibit strong mutual correlations, as demonstrated by the Spearman correlation coefficients (Spearman, 1961) in Figure 3. A strong correlation among the fear-related time series indicates that, regardless of the specific source of information, the perceived fear tends to vary consistently across sources. A lower correlation is observed for the two outliers mentioned above, as well as for None of the above, which exhibits a flatter trend. However, these three time series are highly correlated with one another and with Journalists, indicating that individuals relying on personal or alternative information sources may develop perceptions of fear that differ from those shaped by institutional or official sources. This may point to the existence of an alternative information ecosystem—insulated from broader media trends—where emotional responses are shaped by interpersonal communication or internal beliefs rather than external news flows.
Figure 4. Boxplots showing the distribution of the average fear score across the nine singleton information sources, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups.
Finally, Figure 4 displays boxplots illustrating the distribution of the average fear score across the nine singleton information sources, stratified by Age and Education groups. By excluding the temporal dimension, these distributions offer a concise statistical summary of how fear scores vary across different age and education strata. Additionally, it includes the overall distribution without stratification by Age or Education—note that the overall boxplot is not merely the average of the stratified ones, as the estimates of $\phi_{s^{\prime}}$ were obtained using separate linear models for each grouping variable, as detailed in Equation 5. Specifically, focusing on Age, the average fear score increases with age, with $Age_{1}$ exhibiting the lowest fear score and $Age_{3}$ the highest, across almost all sources—a few exceptions to this trend are observed for CDC, Doctors, and Scientists, where the pattern differs slightly. When focusing on Education, the average fear score generally decreases from $Edu_{1}$ to $Edu_{2}$. In most cases, the fear score then increases again with $Edu_{3}$, although there are some exceptions where it continues to decrease. This behavior highlights the nuanced effect of education on perceived fear, suggesting that while higher education may generally be associated with greater awareness and concern, its influence can vary depending on the source of information and potentially other demographic or contextual factors. In addition, Religious leaders and Family and friends continue to stand out as outliers in the distribution.
5. Causal inference model
The findings from the previous section indicate that the fear score is strongly correlated not only with the sources from which people obtain their information but also with their age and education. Together, these factors contribute significantly to variations in fear across demographic groups. Nonetheless, identifying the primary driver of perceived fear remains challenging. Therefore, we employ causal inference methods to investigate these factors and determine which variable among Source, Age, and Education has the strongest influence on Fear.
Specifically, causal inference involves identifying cause-and-effect relationships by distinguishing them from correlations. It employs statistical methods and study designs—such as randomized controlled trials and quasi-experimental approaches—to estimate the impact of an exposure or intervention on an outcome, while also accounting for confounding factors. While randomized controlled trials represent the gold standard for establishing causality, they are often impractical or unethical. In such cases, causal inference methods can be applied to observational data, provided certain conditions are met—for example, that all relevant variables influencing both the treatment and the outcome are observed, and that the effect of the treatment is consistent across individuals. Common approaches include propensity score matching, instrumental variable analysis, and structural causal models (Pearl, 2010).
Figure 5. Causal inference model used.
Figure 5 illustrates the causal inference model employed in this study, involving four variables: Fear, Source, Age, and Education. Specifically: i) Source $\rightarrow$ Fear denotes the direct causal effect of information source on the outcome variable, the fear score; ii) Age $\rightarrow$ Fear and Education $\rightarrow$ Fear indicate a direct causal pathway from demographic factors to Fear, capturing age- and education-related differences in Fear independently of other influences; iii) Age $\rightarrow$ Source and Education $\rightarrow$ Source represent confounding paths where demographic factors influence the distribution or assignment of individuals to different information sources, potentially confounding the effect of Source on Fear; and iv) Age $\rightarrow$ Education reflects the causal effect of Age on Education level. Age, Education, and Source are categorical variables with Age, Education $\in$ {1, 2, 3} and Source $\in$ {1, 2, …, 9}—each representing one of the nine singleton information sources—, while Fear is a continuous variable.
To define and analyze the causal inference model, we employed the DoWhy package (Sharma and Kiciman, 2020; Blöbaum et al., 2024), a Python library that combines graphical models and statistical methods to specify assumptions, identify causal effects, and estimate them using various techniques. It also provides tools for robustness testing and counterfactual analysis, enabling rigorous causal inference from observational data.
We systematically estimate the strength of each direct causal pathway leading to Fear. This approach allows us to quantify the individual contribution of each variable to variations in Fear, providing a clearer understanding of the underlying causal structure. Our analysis reveals that Source emerges as the most influential variable, explaining a substantial 73.8% of the variance in Fear. This indicates that this variable plays a dominant role in shaping fear responses within the population studied. In comparison, demographic factors such as Age and Education contribute more modestly, accounting for 6.9% and 6.6% of the variance, respectively. These findings suggest that while individual characteristics do have some impact, the Source variable is the primary driver. The remaining 12.7% of variance is attributed to other factors or unobserved variables not included in our current model, highlighting the potential influence of additional elements beyond those considered. This comprehensive estimation underscores the value of causal inference methods in disentangling complex relationships within observational data.
6. Geographical distribution of sources
Figure 6. Temporal evolution of the singleton information sources $\hat{\sigma}_{s^{\prime}}$—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
U.S. nationwide. Using the definitions introduced in Section 3, we analyze patterns in the use of each singleton information source, independently of Fear, represented by $\hat{\sigma}_{s^{\prime}}$. As before, we begin by analyzing the nine distributions; the Shapiro–Wilk test indicates that none follow a normal distribution. To assess differences—including both intra-group and inter-group comparisons—we used the Kolmogorov–Smirnov and Mann–Whitney U tests. The observed significant differences reinforce the idea that perceived uncertainty varies meaningfully across demographic segments and source groupings.
Based on this evidence, Figure 6 shows the evolution of the fraction of respondents relying on each information source, alongside the official number of infections across the U.S. It is notable that None of the above consistently shows values near zero with minimal variation, appearing as an outlier; therefore, we chose not to include it in the visualization. This suggests that very few survey respondents rely on information sources outside those listed in the question. Similarly to the fear score, the reliance of respondents on information sources closely tracks infection trends, exhibiting strong correlation—as shown in Figure 7—and increased sensitivity to relative increases in case numbers rather than absolute values—with the notable exception of Politicians, whose pattern remains flatter and largely uncorrelated. Some sources exhibit moderate sensitivity. The coefficient of variation varies between approximately $0.25$ for None of the above to around $0.1$ for Doctors and CDC, and is lower for other sources, indicating moderate to low variability.
Figure 7. Spearman correlation coefficients among the time series shown in Figure 6.
Respondents’ use of information sources displays strong mutual correlations, as shown in Figure 7, which presents the Spearman correlation coefficients. Within this context, three sources emerge as outliers: Religious leaders—as previously noted—, None of the above, and Politicians. Specifically, Religious leaders and None of the above display negative correlations with all other time series and a positive correlation with each other. This pattern suggests that their trends move in the opposite direction relative to the other sources, indicating that as infection rises, individuals tend to shift away from these less conventional or less trusted sources toward more reliable information channels—an inference supported by their negative correlation with official infection numbers. In contrast, Politicians exhibit negative correlations with Religious leaders, None of the above, and Doctors, while showing a small positive correlation with the other sources. As illustrated in Figure 6, the time series for Politicians remains relatively flat, implying that public reliance on political figures as information providers remains largely stable over time, despite fluctuations in infection severity.
Figure 8. Boxplots showing the distribution of the fraction of respondents relying on each information source, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
Finally, Figure 8 displays boxplots illustrating the distribution of the fraction of individuals relying on each singleton information source, stratified by Age and Education groups, alongside the overall distribution aggregated without stratification. In this case as well, we chose not to include None of the above in the visualization. In both instances, it stands out as a clear outlier. Three clear patterns emerge when moving from $Age_{1}$ to $Age_{3}$ and from $Edu_{1}$ to $Edu_{3}$. First, certain sources maintain a relatively stable or flat level of reliance across these groups, suggesting that demographic factors have little effect on their usage—like Family and friends, Politicians, and None of the above. Second, Religious leaders exhibit a decreasing trend, with younger or less educated groups relying on them more than older or more educated ones. Third, the remaining sources—institutional or expert ones—show increased reliance as Age and Education rise—meaning older or more educated individuals tend to trust these sources more.
U.S. states. To further understand patterns in the use of information sources, we conduct a geographical analysis across U.S. states. The United States presents a diverse socio-economic landscape, with substantial variation in income levels, educational attainment, internet access, and cultural norms across regions. These differences are likely to influence how individuals seek and trust information, particularly in the context of a public health crisis such as the COVID-19 pandemic. By examining the spatial distribution of preferred information sources, we aim to uncover regional trends and disparities that may reflect underlying socio-demographic and economic characteristics. Hence, for each U.S. state, we compute the average reliance on each singleton information source—from $\hat{\sigma}_{s^{\prime}}$—, without stratification by Age or Education, resulting in nine values per state. These averages are then used to cluster the states using K-means (Hartigan et al., 1979), a fast and widely used clustering algorithm that partitions data into a predefined number of clusters by minimizing the distance to cluster centers—it assumes spherical, equally sized clusters. Clustering quality is often evaluated using the Silhouette score (Shahapure and Nicholas, 2020), which measures how well points fit within their assigned clusters, with higher values indicating better-defined clusters.
Figure 9. Clusters derived from K-means analysis based on average reliance on each information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, without stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections. Misclassified states are overlaid with a contrasting grid. States with a purple border indicate swing states between the 2020 and 2024 U.S. political elections.
Based on the Silhouette score, we determined that two clusters provide the optimal partitioning. Figure 9 illustrates the resulting clusters, with each state colored according to its assigned cluster. Interestingly, by chance, we obtained two clusters that correspond closely to the two major U.S. political affiliations: Democratic and Republican. Specifically, Figure 9.a) compares the derived clusters to the 2020 U.S. political elections results, while Figure 9.b) shows the comparison with the 2024 elections outcomes. States that were misclassified—i.e., assigned to the cluster contrary to their political affiliation in the given year—are marked with a grid overlay in the contrasting cluster color. We observe that only two states—along with Alaska, which is not shown on the map—were assigned to the opposite cluster in the 2020 elections, while four states—again including Alaska—were assigned to the opposite cluster in the 2024 elections. Interestingly, these states—except for Alaska—are swing states between the 2020 and 2024 political elections, which makes them inherently more difficult to classify. These results suggest that the clustering based on information source reliance closely reflects established political divisions, with misclassifications primarily occurring in politically competitive swing states. This underscores the sensitivity of information consumption patterns to shifting political landscapes.
7. Conclusion
Building on a foundation of behavioral and demographic data, this study explores the factors driving fear of COVID-19 infection, with a specific focus on how information sources and demographic characteristics influence this fear. We examine how various channels—such as social media, traditional journalism, government briefings, and interpersonal communication—are associated with differences in perceived fear across the U.S. population. By analyzing data spanning May 2021 to June 2022, encompassing the Delta and Omicron waves, we capture a critical period of the pandemic marked by significant informational and emotional turbulence.
Recognizing the potential biases inherent in digital platforms—particularly their skew toward users with consistent internet access—we contextualize our findings within the socio-economic and geographic diversity of the United States. Regional disparities in media habits, internet penetration, and education levels emphasize the importance of a geographically nuanced approach to understanding informational dynamics and public emotion.
Our analysis reveals that the type of information source individuals rely on plays a central role in shaping emotional responses during health crises. The credibility, framing, and volume of content can either amplify or mitigate public fear, with misinformation posing a significant threat to risk perception and trust. This underscores the critical importance of ensuring access to verified, reliable information sources to foster informed decision-making and emotional resilience.
Using causal inference techniques, we assessed the relative influence of age, education, and information sources on perceived fear. Among these factors, the type of information source emerged as the most significant driver. Furthermore, our findings demonstrate that demographic characteristics not only influence emotional reactions but also shape preferences for specific types of information, revealing distinct behavioral patterns across age and education groups. Taken together, these insights provide critical guidance for designing more targeted and trustworthy communication strategies in future public health emergencies, where managing both the transmission of disease and the spread of fear is essential.
Moreover, our findings reveal a strong connection between information ecosystems and political behavior. Clustering analysis showed that patterns of reliance on specific information sources across U.S. states closely aligned with political orientations during the 2020 and 2024 election cycles. These results underscore the broader societal impact of informational environments, demonstrating their influence not only on emotional responses but also on civic attitudes and decision-making. In this context, it becomes imperative that official information channels maintain the highest standards of accuracy and credibility. The dissemination of misinformation—whether intentional or due to negligence—by institutional sources not only distorts public understanding but also undermines democratic processes by shaping political beliefs on falsehoods. Therefore, safeguarding trusted information sources from distortion is crucial for protecting both public health and the integrity of democratic society.
Appendix A Extended analysis on clustering
In addition to the analysis presented in Section 6—specifically in the U.S. states subsection—, we provide additional details on clustering using $\hat{\sigma}_{s^{\prime}}$. We also repeat the analysis with alternative inputs, including $\phi_{s^{\prime}}$ alone and in combination with $\hat{\sigma}_{s^{\prime}}$.
Figure 10. Clusters derived from K-means analysis based on average reliance on each singleton information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
Clustering using $\hat{\sigma}_{s^{\prime}}$. Alongside the results shown in Section 6, we incorporate demographic information by computing, for each U.S. state, the average reliance on each singleton information source, grouped by Age, resulting in 27 values per state. As shown in Figure 10, this adjustment reduced misclassifications in the 2024 elections, with only Wisconsin and Alaska assigned to the opposite cluster—but led to an increase in misclassifications for the 2020 elections. A similar improvement is observed when combining information source preferences with Education levels.
The improved classification performance observed for the 2024 elections when incorporating age and education suggests that these demographic variables became more closely associated with political orientation, likely due to shifts in how different age and education groups access and respond to information. In contrast, the reduced accuracy for the 2020 elections indicates that these factors were less informative during that period, possibly due to more uniform voting patterns across demographics or the influence of external events such as the COVID-19 pandemic.
Clustering using $\phi_{s^{\prime}}$. For each U.S. state, we compute the average fear score for each singleton information source, $\phi_{s^{\prime}}$, without stratifying by Age or Education, resulting in nine values per state. Figure 11 presents the resulting clusters and compares them with the outcomes of the 2020 and 2024 political elections.
Compared to the clusters obtained using $\hat{\sigma}_{s^{\prime}}$, a larger number of states are misclassified in both elections—specifically, Colorado, Connecticut, Delaware, Illinois, Maine, Massachusetts, Minnesota, New Hampshire, Oregon, Rhode Island, Texas, Vermont, Virginia, and Washington. Notably, swing states are misclassified only in the 2020 elections, but appear in the correct cluster for 2024. Additionally, Alaska and Hawaii are now correctly classified in both elections.
Next, we stratify the data by Age and Education, yielding 27 values per state. Stratifying by Age yields similar results, with the same set of states appearing misclassified. When stratifying by Education, two additional states—New York and Maryland—are misclassified in both the 2020 and 2024 elections. This shift suggests that, for these states, educational differences may influence the relationship between fear scores and political alignment in a manner that diverges from the general pattern observed without stratification by education.
Figure 11. Clusters derived from K-means analysis based on the average respondents’ fear score for each singleton information source, $\phi_{s^{\prime}}$, across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
More broadly, the presence of traditionally Democratic-leaning states among those misclassified indicates that ideological or cultural factors—such as attitudes toward information sources, regional media ecosystems, or local political discourse—might play a role in shaping fear perception patterns that do not align neatly with national political divisions.
Clustering using $\hat{\sigma}_{s^{\prime}}$ and $\phi_{s^{\prime}}$. For each U.S. state, we compute the average reliance $\hat{\sigma}_{s^{\prime}}$ and fear scores $\phi_{s^{\prime}}$ for each singleton information source without stratifying by Age or Education, resulting in 18 values per state. Figure 12 presents the resulting clusters obtained and compares them to the outcomes of the 2020 and 2024 U.S. political elections.
When combining both $\hat{\sigma}_{s^{\prime}}$ and $\phi_{s^{\prime}}$, we observe that five states—Alaska, Georgia, Michigan, Pennsylvania, and Wisconsin—are misclassified for the 2020 elections, while only three states—Alaska, Arizona, and Nevada—are misclassified for the 2024 elections.
Next, we stratify the data separately by Age and Education for both $\hat{\sigma}_{{s^{\prime}}}$ and $\phi_{{s^{\prime}}}$, yielding 54 values per state. Stratifying by Education yields similar results, with the same set of states appearing misclassified. In contrast, when stratifying by Age, only Alaska remains in the wrong cluster for the 2024 elections. This suggests that age-related differences in information consumption and fear perception play a key role in accurately classifying Arizona and Nevada. In these states, incorporating age stratification captures important nuances in how fear and reliance on information sources vary across generations, ultimately aligning their political clustering more accurately.
Interestingly, Wisconsin exhibits the opposite pattern in the 2024 elections. While clustering with $\hat{\sigma}_{s^{\prime}}$ and stratifying by age alone is insufficient for correct classification, incorporating fear scores—or a combination of fear and reliance for each information source—is essential for accurate placement. This suggests that, in Wisconsin, the emotional response to information sources plays a more decisive role in shaping political alignment than demographic factors alone.
Figure 12. Clusters derived from K-means analysis based on the average reliance $\hat{\sigma}_{s^{\prime}}$ on each singleton information source and average respondents’ fear score $\phi_{s^{\prime}}$ for them across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
Appendix B Ethical declaration
The Ethics Board (IRB) of IMDEA Networks Institute gave ethical approval for this work on 2021/07/05. IMDEA Networks has signed Data Use Agreements with Facebook, Carnegie Mellon University (CMU) and the University of Maryland (UMD) to access their data, specifically UMD project 1587016-3 entitled C-SPEC: Symptom Survey: COVID-19 and CMU project STUDY2020_00000162 entitled ILI Community-Surveillance Study. The data used in this study was collected by the University of Maryland through The University of Maryland Social Data Science Center Global COVID-19 Trends and Impact Survey in partnership with Facebook. Informed consent has been obtained from all participants in this survey by this institution. All the methods in this study have been carried out in accordance with relevant ethical and privacy guidelines and regulations.
Appendix C Availability of materials and data
All datasets used in this work are open-source and publicly available. In particular, we used official surveillance data from the COVID-19 Data Hub (Guidotti and Ardia, 2020; Guidotti, 2022) and additional data from the Delphi US CTIS (Delphi Group, CMU, 2020) on demographic characteristics (such as age and education), sources of COVID-19-related information, and reported levels of fear. The microdata of the Delphi US CTIS survey from which the aggregated data was obtained cannot be shared, as per the Data Use Agreements signed with Facebook, Carnegie Mellon University (CMU), and the University of Maryland (UMD). Thus, the authors do not have permission to share the microdata of the Delphi US CTIS survey. The Delphi US CTIS survey data (in aggregated form) and the code are openly accessible at https://github.com/paolocastagno/covid19-fear.
Appendix D Declaration of competing interest
The authors declare no competing interests.
Appendix E Funding
This work was supported by grants from TED2021-131264B-I00 (SocialProbing) funded by MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR.