### v2

# Whom We Trust, What We Fear: COVID-19 Fear and the Politics of Information

Abstract
The COVID-19 pandemic triggered not only a global health crisis but also an infodemic, an overload of information from diverse sources influencing public perception and emotional responses. In this context, fear emerged as a central emotional reaction, shaped by both media exposure and demographic factors. In this study, we analyzed the relationship between individuals’ self-reported levels of fear about COVID-19 and the information sources they rely on, across nine source categories, including medical experts, government institutions, media, and personal networks. In particular, we defined a score that ranks fear levels based on self-reported concerns about the pandemic, collected through the Delphi CTIS survey in the United States between May 2021 and June 2022. We found that both fear levels and information source usage closely follow COVID-19 infection trends, exhibit strong correlations within each group (fear levels across sources are strongly correlated, as are patterns of source usage), and vary significantly across demographic groups, particularly by age and education. Applying causal inference methods, we found that among age, education, and information source, the latter is the most influential factor affecting individuals’ fear levels. We further quantified the impact of different information sources on fear by estimating the average treatment effect, indicating how each source alters fear relative to a control. Furthermore, we demonstrated that information source preferences can reliably match the political orientation of U.S. states. These findings highlight the importance of information ecosystem dynamics in shaping emotional and behavioral responses during large-scale crises.
Introduction
The COVID-19 pandemic represented a global health crisis of unprecedented scale and complexity. At the same time, it generated an infodemic—a rapid and overwhelming surge of information from a wide range of sources (Cinelli and others 2020). As individuals sought clarity and guidance, they turned to news media, social platforms, government announcements, and public health organizations. The credibility, consistency, and framing of these sources played a critical role in shaping public understanding of the threat. Additionally, they influenced emotional responses across populations (Jones et al. 2021; Lu and others 2021).
In this context, fear emerged as one of the most prominent psychological responses to the pandemic (Coelho et al. 2020). This emotion was not driven solely by the objective health threat, but also by the volume, tone, and perceived trustworthiness of the information consumed. The impact of these information sources on fear varied across different demographics. Differences in age, education, gender, and socioeconomic status significantly shaped how individuals selected, interpreted, and trusted various sources (Nino et al. 2021). These disparities highlight the necessity of accounting for both the nature of the information disseminated and the characteristics of the recipients when evaluating the psychological impact of information exposure.
Furthermore, information played a dual role during the crisis: it served both as a practical resource for guiding health-related decisions and as a symbolic indicator of political identity and civic orientation. The choice to trust or dismiss specific sources frequently mirrored broader ideological affiliations, shaping not only individual behaviors but also the level of collective trust in public institutions. Against this backdrop, the risks posed by misinformation—particularly when it stemmed from institutional or otherwise authoritative sources—became especially salient. Perceptions of inconsistency, politicization, or unreliability in official communications had the potential to undermine not only adherence to public health guidance but also confidence in democratic institutions. Understanding how various demographic groups navigated this complex informational environment is, therefore, critical for designing communication strategies that bolster both public health outcomes and civic trust in future crises.
Figure 1: Schematic representation of our study.
Contributions
Our main contributions are as follows—summarized in Figure 1. First, using statistical tests and correlation analyses on U.S. data, we quantify the influence of different information sources on individuals’ COVID-19–related fear levels, and similarly assess differences in information source usage across the population. Second, we investigate how demographic factors—specifically age and education—not only affect emotional responses but also significantly shape information-seeking behavior and trust, revealing notable patterns across population groups. Third, through a causal inference model, we compare the relative impact of age, education, and information source on fear, finding that information source is the most influential driver. We then estimate the average treatment effect (ATE) to quantify how each source alters fear compared with a control source. Finally, building on this, we apply K-means, spectral, and hierarchical clustering to group U.S. states based on their residents’ reliance on different types of information. Remarkably, the resulting clusters exhibit a strong alignment with state-level political orientation—specifically, whether a state tends to lean Democratic or Republican. Collectively, our findings underscore the critical role of information sources in shaping population emotional responses to large-scale crises. They also highlight the dangers of misinformation—particularly when disseminated by institutional or authoritative entities—which can distort public perception, exacerbate fear, and undermine trust in democratic institutions.
This article is organized as follows. The Related work section reviews prior research on fake news, misinformation, and COVID-19–related fear, highlighting key findings and existing gaps. The Data section describes our data sources and the construction of the indices used in the analysis. The Methods section presents the methodology, including correlation analysis, causal inference, and clustering. The Results section examines variation in fear levels across information sources and demographic groups, as well as variation in information source usage across demographics, and identifies patterns of information reliance across the U.S. Finally, the Conclusion section summarizes the main findings and discusses their broader implications.
Related work
A growing body of research has examined the spread of fake news and misinformation during the COVID-19 pandemic and their psychological consequences—particularly the prevalence and dynamics of fear. This fear has been fueled by false or low-credibility information presented as legitimate news which has also undermined trust in vaccines and preventive measures and contributed to a global infodemic (Caceres and others 2022).
Evidence shows that information sources strongly shape risk perceptions and emotional responses. In Portugal, scientific sources had greater effects on perceived personal and societal risk than general media, and trust in science heightened perceived vulnerability, while trust in social media reduced it (Entradas 2022). In the U.S., socioeconomic and political divisions shaped information-seeking behaviors: individuals with higher education and digital skills accessed a more diverse range of sources, whereas reliance on conservative media was associated with lower levels of factual knowledge (Reisdorf and others 2021). Additional analyses of social media behavior show that online discourse around COVID-19 was heavily politicized: an investigation of Facebook posts from users who engaged with anti-government and COVID-denialist narratives before later dying from the virus found that harmful themes were largely propagated within a right-wing political and media ecosystem, with platforms applying only soft moderation measures (Habib and Nithyanand 2023). This underscores both the politicization of COVID-19 information and the role of platform governance in shaping exposure to dangerous narratives.
False information spreads far more rapidly than accurate content on social platforms, largely propagated by superspreaders in ideologically aligned networks. While fact-checking demonstrates limited durability, psychological inoculation strategies employing gamified prebunking show enhanced efficacy in building cognitive resistance (Van Der Linden 2022). These dynamics amplify infectious disease transmission through behavioral mechanisms by promoting harmful practices and undermining prevention. Misinformation persists longer and spreads more readily than factual information, substantially increasing infection rates (Bernardin and Perez-Acle 2025). Additionally, the impact on vaccination intentions is especially pronounced; randomized controlled trials demonstrate that exposure to vaccine misinformation reduces the intent to vaccinate by approximately 6 percentage points relative to factual information, with greater effects for scientifically sounding misinformation and differential impacts across sociodemographic groups (Loomba and others 2021). Research also shows that misinformation about COVID-19 vaccines is driven by coordinated conspiracy-oriented communities that amplify anti-vaccine narratives and distort scientific evidence, contributing to polarized uptake patterns (Sharma et al. 2022). Automated approaches to misinformation monitoring also support these findings: neural stance-classification models using graph attention networks can reliably identify whether social media posts adopt or reject COVID-19 misinformation, revealing consistent patterns in the emotional, lexical, and thematic structure of misinformation uptake (Weinzierl et al. 2021). Meanwhile, even a single trusted correction can reduce belief in misinformation (Seo et al. 2022; Miyazaki and others 2023), emphasizing the importance of timely and credible debunking.
Social media commentary on COVID-19 was largely negative, with anger dominating globally and fear peaking early in the pandemic, later rising only during variant or mortality news. Emotional expression varied by region, with the U.S. showing more anger, African outlets more sadness, and Indian sources more fear. In particular, fear emerges as a central psychological response shaped more by subjective perceptions and social transmission than by epidemiological indicators. Social media analyses reveal that fear expression peaks during early outbreak phases and shows substantial gender differences, with women expressing fear more frequently than men in response to disease-related topics. The temporal dynamics of fear demonstrate marked fluctuations independent of actual disease severity, with daily shares of fear posts significantly elevated during peak COVID-19 periods compared to baseline 2019 levels (Chai and others 2022). Cross-national longitudinal research finds that roughly 23% of fear variance is explained by micro-level psychological factors—worries about resource shortages, perceived disease vulnerability, germ aversion, and infections in close social circles—while macro-level indicators like case counts, mortality, and restrictions contribute minimally. Individual traits such as age, attachment anxiety, and prior health concerns further shape fear responses, revealing heterogeneous susceptibility and a disconnection between subjective threat perception and objective epidemiological data (Eder and others 2021). Agent-based models with associative learning reveal that falling fear can trigger new disease waves by reducing protective behavior, which strongly shapes epidemic trajectories. Media campaigns promoting protective actions have minimal impact, suggesting interventions must target the psychological drivers of fear and compliance, not just factual disease information (Retzlaff and others 2022).
Moreover, early population-level fear modestly predicted later vaccination willingness but was insufficient alone to explain vaccine hesitancy, which involves multiple psychological and risk-perception factors. Demographically, females, healthcare workers, lower-educated, and married individuals showed higher fear, highlighting vulnerable groups needing targeted interventions that address fear alongside vaccine-related beliefs (Oliveira and others 2023; Mertens et al. 2022; Doshi and others 2021). Finally, causal inference reveals bidirectional relationships between media exposure, interpersonal communication, and fear of COVID-19. Both media frequency and perceived danger in conversations predict fear, while elevated fear motivates increased media and communication seeking. Counterarguing proves most effective at reducing fear and encourages sustained use, with individuals cycling from avoidance to counterarguing to reappraisal as the pandemic progresses. These adaptive emotion-regulation patterns suggest that targeted interventions addressing maladaptive regulatory pathways prove more effective than uniform communication approaches (Li 2022).
Overall, misinformation and fear jointly influence pandemic behavior through interconnected mechanisms. Misinformation spreads rapidly, reduces vaccination intentions, and undermines protective behaviors, while fear—driven largely by subjective perceptions, information exposure, and social context—varies across demographic groups and feeds back into information-seeking patterns. Declining fear can weaken compliance and contribute to renewed outbreaks.
Question Possible responses
(D2) What is your age? (1) 18-34 years, (2) 35-64 years, (3) 65+ years
(D8) What is the highest degree or level of school you have completed? (1) Lower than high school, (2) High school / Bachelor’s degree, (3) Postgraduate
(I5) In the past 7 days, from which of the following sources have you received news and information about COVID-19? (1) Doctors, (2) Scientists, (3) CDC, (4) Government, (5) Politicians, (6) Journalists, (7) Family and friends, (8) Religious leaders, (9) None of the above (other sources not listed here).
(G1) How much do you worry about catching COVID-19? (1) A great deal, (2) A moderate amount, (3) A little, (4) Not at all.
Table 1: Selected questions and responses from the Delphi US CTIS survey, with labels as in the original survey.
Despite extensive research on misinformation and fear during the COVID-19 pandemic, key gaps remain. The influence of different information sources on fear across demographic groups remains poorly characterized. Additionally, geographic and political influences remain underexplored. Addressing these gaps is critical for designing interventions that effectively mitigate misinformation, target heterogeneous fear responses, and support vulnerable populations. Our study addresses these gaps by employing advanced statistical and causal inference methods to evaluate how demographics and sources of information jointly shape fear responses. Moreover, by clustering U.S. states according to patterns of information source usage and linking these clusters to political orientation, we offer a comprehensive, integrated perspective on the social and contextual determinants of COVID-19 fear.
Data
Throughout this work, we relied on two sources of COVID-19 data: official surveillance data from the COVID-19 Data Hub (Guidotti and Ardia 2020; Guidotti 2022) and survey-based data from the Delphi US CTIS (Delphi Group, CMU 2020), which contains information on demographics, the types of information sources used by individuals, and self-reported levels of fear. All datasets used in this study are open-source and publicly available. Aggregated data from the Delphi US CTIS survey can be accessed as described in the Availability of materials and data section.
We primarily used the COVID-19 Data Hub (Guidotti and Ardia 2020; Guidotti 2022) as our source of surveillance data. Specifically, we used daily new fatalities and confirmed cases—computed from cumulative counts—as well as the total U.S. population. These indicators were used to estimate the daily number of official infections via Sybil (Baccega et al. 2024). Specifically, this data allows us to reconstruct the dynamics of a Susceptible-Infected-Recovered-Deceased-Susceptible (SIRDS) compartmental model with reinfections over the study period. In this model, the infected compartment represents the estimate of the official number of infections (Álvarez and others 2021; Astley and others 2021)—we assumed an average recovery time of 14 days to compensate for the lack of recovery data, which are largely unavailable. While the Data Hub provides additional information, such as vaccination rates and hospital statistics, these were not relevant to our study.
Since the spring of 2020, Facebook has collaborated with both Carnegie Mellon University (CMU) and the University of Maryland (UMD) through its Data for Good program (Facebook 2020) to promote large-scale, daily surveys developed by these institutions (Astley and others 2021; Fan and others 2020). The Delphi Group at CMU administers the U.S. COVID-19 Trends and Impact Survey (Delphi Group, CMU 2020)—Delphi US CTIS—, which collects approximately 50,000 responses per day from Facebook users in the United States. This survey includes self-reported information on symptoms, testing, vaccination, behaviors, and other COVID-19-related topics, and was used to estimate prevalence trends across U.S. states, accounting for direct and indirect reporting (Srivastava et al. 2024). Access to the data was obtained through an agreement with CMU, UMD, and Facebook, as described in the Ethical Declaration section. The full set of survey questions in the U.S. is available at (Delphi Group, Carnegie Mellon University 2020).
Specifically, we computed daily aggregated information considering questions about age, education, source of information, and infection-induced fear—a summary of the selected questions and their corresponding response options is provided in Table 1. Importantly, these daily aggregates were calculated using the weights provided in the Delphi US CTIS microdata files. These weights were designed to address the differences between Facebook users and the overall US population, as well as to correct for biases associated with non-response (Salomon and others 2021). For clarity and simplicity, throughout the remainder of the paper, we refer to responses to D2 as Age, D8 as Education, I5 as Source, and G1 as Fear. For Source, users could select multiple options, whereas for Age, Education, and Fear responses are mutually exclusive. Therefore, we accounted for all possible combinations of the nine response options in Source, totaling 256—also comprising the None of the above response, which is used in mutual exclusion with all the other possible answers. We will focus on the relationships between Age, Education, and Source with respect to COVID-19–induced Fear, employing the notation presented in Table 2.
Our analysis covers the period from May 2021 to June 2022 across all U.S. states, which combines online activity with self-reported information. While these data sources are inherently biased toward individuals with consistent internet access—specifically on active Facebook users—and may underrepresent certain socio-demographic groups or regions with limited connectivity, these biases are less pronounced in U.S. states where internet access is widespread—hence, offering valuable insights into broader behavioral trends.
Normalization. This step primarily facilitates the later examination of correlations and comparisons among variables. By grouping respondents based on Age, Education, or not grouping at all—denoted by variable vv—, it is straightforward to derive the normalized quantities $\hat{\sigma}_{s}$:
$\hat{\sigma}_{s}(v)=\frac{\sigma_{s}(v)}{\sum_{i\in\mathcal{S}}\sigma_{i}(v)},$ (1)
where we dropped the time dependence for ease of notation. Similarly, we define the normalized Fear measure $\hat{\varphi}_{s,f}$ for different levels of self-reported fear, applying an analogous procedure.
Symbol Description
$\mathcal{S}$ Set of 256 combinations of Source [^1] In detail, $\mathcal{S}=\big\{${Doctors}, {Scientists}, {CDC}, $\dots$, {None of the above}, {Doctors, Scientists}, {Doctors, CDC}, $\dots\big\}$ is a set containing 256 subsets.
$s^{\prime}$ Singleton information sources from $\mathcal{S}$, e.g., $\{\text{{Doctors}}\}$ and $\{\text{{Scientists}}\}$
$\mathcal{F}$ Set of self-reported fear levels $f\in\mathcal{F}$, where $\mathcal{F}=\left\{1,\ 2,\ 3,\ 4\right\}$
$\sigma_{s}$ Average number of respondents selecting the combination $s\in\mathcal{S}$
$\sigma_{s^{\prime}}$ Average number of respondents selecting singleton $s^{\prime}\in\mathcal{S}$
$\varphi_{s,\ f}$ Average number of respondents who self-reported a fear level $f\in\mathcal{F}$, selecting the combination $s\in\mathcal{S}$
$\varphi_{s}$ Average fear score among respondents selecting the combination $s\in\mathcal{S}$
$\phi_{s^{\prime}}$ Estimate of average fear score for singletons $s^{\prime}\in\mathcal{S}$
$v$ Demographic grouping: Age, Education, or no grouping
$t$ Time variable $t\in\left[1,\ T\right]$, where $T$ denotes the length of the time series, with one observation per day
Table 2: Notation used in the paper.
Singleton information sources. Data on information usage is provided in aggregated form, where each $s\in\mathcal{S}$ represents a possible combination of the considered information sources. However, to investigate the utilization of individual sources, we extrapolate the proportion of respondents using each singleton information source. Specifically, the average number of people using a specific singleton source $\hat{\sigma}_{s^{\prime}}$ is given by
$\hat{\sigma}_{s^{\prime}}(v)=\sum_{\forall u\in\mathcal{S}:s^{\prime}\in u}\hat{\sigma}_{u}(v),$ (2)
where $s^{\prime}$ are the singletons information sources from $\mathcal{S}$. Please note that, although $\hat{\sigma}_{s^{\prime}}$ is computed starting from the normalized variable $\hat{\sigma}_{s}$, the sum of $\hat{\sigma}_{s^{\prime}}$ over all singletons $s^{\prime}$ might exceed 1 since individual values of $\hat{\sigma}_{u}$ are counted once for each $s^{\prime}\in u$.
Fear score. By assigning weights $w_{f}$ to each fear level, we define the fear score as
$\hat{\varphi}_{s}(v)=\sum_{f\in\mathcal{F}}w_{f}\cdot\hat{\varphi}_{s,\ f}(v),$ (3)
where $w_{f}\in\{1,\ 0.34,\ -0.34,\ -1\}$. This formulation ensures that responses such as A great deal are counterbalanced by Not at all, with a similar compensatory relationship between A moderate amount and A little.
Fear score of singleton information sources. In case of Fear, we cannot directly apply the approach used in Equation 2 to obtain the average fear score from singleton source responses. This limitation arises because, for each $s\in\mathcal{S}$, the distribution of Fear across the selected sources is unknown. For example, $s=\{$CDC, Politicians, Journalists}\} denotes the average fear score across all U.S. states among respondents who selected CDC, Politicians, and Journalists as their information sources. However, this aggregated value does not reveal how much each source contributes to the self-reported fear—this could be primarily driven by Journalists, equally influenced by all three sources, or dominated by a different source within the combination.
To address this, and estimate $\phi_{s^{\prime}}$, we employ linear models to disentangle the contribution of each source to the overall average fear score as
$\min_{w_{s^{\prime}}|s}\sum_{t\in T}\left(\hat{\varphi}_{s}(v,\ t)-\sum_{s^{\prime}\in s}w_{s^{\prime}|s}\cdot\hat{\varphi}_{s^{\prime}}(v,\ t)\right)^{2},$ (4)
where $w_{s^{\prime}}|s$ represents time-independent weights assigned to each singleton source $s^{\prime}$ in the combination ss—this optimization problem is solved for each $s\in S$. These weights are estimated by minimizing the squared difference between the observed average fear score of the combination $s$, denoted $\hat{\varphi}_{s}(v,\ t)$, and the weighted sum of the average fear scores of its singleton sources, $\hat{\varphi}_{s^{\prime}}(v,\ t)$. Specifically, we use $\hat{\varphi}_{s^{\prime}}(v,\ t)$—the fear score among respondents who selected $s^{\prime}$—as a proxy to estimate the weights. In this way, for each $\hat{\varphi}_{s}(v)$, we determine the contribution of each singleton $s^{\prime}\in s$. Using these weights, we then compute $\phi_{s^{\prime}}$ as
$\phi_{s^{\prime}}(v)=\sum_{s\in\mathcal{S}}w_{{s^{\prime}}|s}\cdot\hat{\varphi}_{s^{\prime}}(v)$ (5)
Only weights from models with statistically significant coefficients—determined via F-tests—are retained to ensure robust attribution.
Methods
This section describes the primary methodology used in our study, as illustrated in Figure 1.
Statistical analysis on fear. Using U.S. data, we assess how different information sources affect COVID-19–related fear through statistical tests—Shapiro–Wilk (Shapiro and Wilk 1965), Kolmogorov–Smirnov (Massey Jr 1951) and Mann–Whitney U (Nachar and others 2008) tests—and Spearman correlation analyses (Spearman 1961). We further examine how demographic factors—specifically age and education—shape emotional responses, thereby revealing distinct patterns across population groups. Specifically, the Shapiro–Wilk test was used to assess whether a distribution follows normality, whereas the Kolmogorov–Smirnov test was applied to compare two distributions, evaluating differences across their full range, including shape, dispersion, and central tendency. In contrast, the Mann–Whitney U test was employed to detect differences in central tendency between two distributions. The Spearman correlation coefficient was used to quantify the strength and direction of monotonic relationships, allowing us to examine how fear levels co-vary with different information sources.
Causal inference. Using causal inference, we evaluate the relative contributions of age, education, and information source to COVID-19–related fear, finding that information source is the most influential factor. In particular, causal inference aims to uncover cause-and-effect relationships rather than mere statistical associations. It employs a range of statistical methods and study designs—such as randomized controlled trials and quasi-experimental approaches—to estimate the effect of an exposure or intervention on an outcome while accounting for potential confounders. Randomized controlled trials remain the gold standard for establishing causality, as randomization helps balance confounding factors across groups. However, these trials are often infeasible or ethically problematic. In such cases, causal inference methods can be applied to observational data, provided key assumptions hold—consistency, positivity, and exchangeability. Moreover, we estimate the average treatment effect (ATE) of different information sources, where the ATE—one of the primary metrics in causal inference—quantifies the expected difference in outcomes between a control treatment (in our case, None of the above) and each alternative information source, thereby reflecting the causal impact of the intervention at the population level.
Figure 2: Causal inference model used.
To define the causal inference model—shown in Figure 2—, we use the DoWhy package (Sharma and Kiciman 2020; Blöbaum and others 2024), a Python library that integrates graphical causal models with statistical methods to specify assumptions, identify estimands, estimate causal effects, and perform robustness checks, including refutation tests. DoWhy was mainly used to estimate the strength of causal relationships (arrow strengths) by fitting a Probabilistic Causal Model (PCM) to the data. For effect estimation, however, we used Ordinary Least Squares (OLS) regression models, adjusting for confounders (Age and Education) identified in the causal graph, to obtain ATE estimates. DoWhy’s refutation API was applied to validate model robustness through tests like placebo treatments and subset analyses, ensuring the reliability and interpretability of our OLS-based causal estimates from observational data.
In detail, the positivity assumption was verified by confirming that all information sources appear across all Age and Education strata, ensuring each source has a positive probability of selection. Each source was treated as a binary treatment (selected or not on a given date), assuming consistent definitions, no meaningful heterogeneity in usage, and additive effects. While the intensity and context of source use may vary (e.g., passive exposure vs. active seeking), such distinctions are not captured in the survey. The analysis further assumes that age and education suffice to control for confounding, though unmeasured factors—such as political ideology, institutional trust, socioeconomic status, and personal traits—may influence results. Consequently, estimated causal effects reflect associations conditional on measured confounders, and causal interpretation relies on the unverifiable assumption of no relevant unmeasured confounding. Sensitivity analyses for unmeasured confounding are beyond the scope of this study.
Geographical distribution of sources. Finally, we statistically assess how people use different information sources using the previously described tests and Spearman correlation analyses, while also examining how demographic factors influence information source usage, thereby revealing distinct patterns across population groups. Next, we apply K-means (Hartigan et al. 1979), spectral (Ng et al. 2001), and hierarchical clustering (Johnson 1967) to group U.S. states based on their residents’ reliance on different types of information. Remarkably, the resulting clusters exhibit a strong alignment with state-level political orientation—specifically, whether a state tends to lean Democratic or Republican. In detail, K-means partitions data into $k$ clusters by iteratively assigning points to the nearest centroid, while spectral clustering uses graph Laplacian eigenvectors to capture non-convex structures before applying a partitioning method such as K-means. Hierarchical clustering builds a dendrogram through agglomerative or divisive steps, allowing flexible cluster counts but at a higher computational cost. Cluster quality is commonly assessed using the Silhouette score, which quantifies how well points fit within their assigned clusters.
Results
Figure 3: Temporal evolution of $\phi_{s^{\prime}}$ for each of the nine singleton information sources—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. To reduce noise, official infections and $\phi_{s^{\prime}}$ were smoothed using a two-month rolling average.
Statistical analysis on fear. Employing the definitions introduced in the Data section, we investigate the drivers behind the evolution of the nationwide U.S. fear score $\phi_{s^{\prime}}$. As a starting point, we examine whether Fear differs across the various information sources. In particular, considering the nine distributions derived from the singleton information sources $s^{\prime}$, we first applied the Shapiro–Wilk test, which indicated that none of the distributions are normally distributed. Next, to assess whether these distributions differ significantly—both within groups (across the nine singleton information sources for each Age or Education group and at the population level) and between groups (across different Age or Education groups)—we applied the Kolmogorov–Smirnov and Mann–Whitney U tests. The consistently significant results indicate meaningful variation in fear-score distributions at both levels, reflecting demographic and informational heterogeneity.
In light of these findings, Figure 3 presents the temporal evolution of the singleton fear scores $\phi_{s^{\prime}}$, alongside the official number of infections across the U.S. Notably, Religious leaders and Family and friends emerge as significant outliers, consistently associated with lower fear scores [^2] Fear score spanning values around zero reflects design choice of assigning weights that counterbalance the associated responses in the survey, as defined in Equation 3.. Moreover, the fear score $\phi_{s^{\prime}}$ closely mirrors fluctuations in official infection numbers, indicating a strong correlation between the two—as shown in Figure 4, where we can notice that $\phi_{s^{\prime}}$ is more sensitive to relative increases in case numbers than to absolute levels.
Figure 4: Spearman correlation coefficients among the time series shown in Figure 3.
Some sources show even greater sensitivity. The coefficient of variation [^3] The coefficient of variation is the standard deviation divided by the mean’s absolute value, measuring relative variability. is particularly high for Scientists ($\approx 3$) and the CDC ($\approx 1$), whereas all other sources remain below $0.5$, indicating moderate variability. This pattern illustrates how fear intensifies with the rising visibility of COVID-19 cases, dominating public discourse. The moderate to high variation highlights its strong dependence on fluctuations in official case numbers.
Among all sources, Family and friends and Journalists slightly anticipate phase transitions in infection trends: their signal $\phi_{s^{\prime}}$ precedes both the initial rise and the peak decline during the Delta and Omicron waves (Dattani and Rodés‑Guirao 2024). This pattern may reflect an anticipatory response driven by early information or heightened awareness within these social and media networks.
In addition, singleton fear scores exhibit strong mutual correlations, as demonstrated by the Spearman correlation coefficients in Figure 4. A strong correlation among the fear-related time series indicates that, regardless of the specific source of information, the perceived fear tends to vary consistently across sources. A lower correlation is observed for the two outliers mentioned above, as well as for None of the above, which exhibits a flatter trend. However, these three time series are highly correlated with one another and with Journalists, indicating that individuals relying on personal or alternative information sources may develop perceptions of fear that differ from those shaped by institutional or official sources. This may point to the existence of an alternative information ecosystem—insulated from broader media trends—where emotional responses are shaped by interpersonal communication or internal beliefs rather than external news flows.
Figure 5: Boxplots showing the distribution of $\phi_{s^{\prime}}$ across the nine singleton information sources, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups.
Finally, Figure 5 displays boxplots illustrating the distribution of the signal $\phi_{s^{\prime}}$ across the nine singleton information sources, stratified by Age and Education groups. By excluding the temporal dimension, these distributions offer a concise statistical summary of how fear scores vary across different age and education strata. Additionally, it includes the overall distribution without stratification by Age or Education—note that the overall boxplot is not merely the average of the stratified ones, as the estimates of $\phi_{s^{\prime}}$ were obtained using separate linear models for each grouping variable, as detailed in Equation 5. Specifically, focusing on Age, $\phi_{s^{\prime}}$ increases with age, with $Age_{1}$ exhibiting the lowest fear score and $Age_{3}$ the highest, across almost all sources—a few exceptions to this trend are observed for CDC, Doctors, and Scientists, where the pattern differs slightly. When focusing on Education, $\phi_{s^{\prime}}$ generally decreases from $Edu_{1}$ to $Edu_{2}$, and in most cases rises again with $Edu_{3}$, although in some instances it continues to decline. This behavior highlights the nuanced effect of education on perceived fear, suggesting that while higher education may generally be associated with greater awareness and concern, its influence can vary depending on the source of information and potentially other demographic or contextual factors. In addition, Religious leaders and Family and friends continue to stand out as outliers in the distribution.
Causal inference. The findings from the previous section indicate that $\phi_{s^{\prime}}$ varies significantly not only with the sources of information people rely on but also with their age and education. Nonetheless, identifying the primary driver of perceived fear remains challenging. Therefore, we employ causal inference methods to investigate these factors and determine which variable among Source, Age, and Education has the strongest influence on Fear.
Figure 6: Estimated direct effects on Fear. Panels a) and b) show variance explained by each variable in the first and second waves, respectively. Panels c) and d) show the influence of each information source.
Figure 2 illustrates the causal inference model, featuring nine binary Source variables (one-hot encoded from the original categorical source), continuous outcome Fear, and categorical confounders Age, Education $\in\{1,2,3\}$. Direct effects flow from each Source $\rightarrow$ Fear and demographics $\textit{Age}{},\textit{Education}{}\rightarrow$ Fear. Critical confounding arises via $\textit{Age}{},\textit{Education}{}\rightarrow$ Source, where demographics determine information source exposure, biasing naive source-fear associations; additionally, $\textit{Age}{}\rightarrow$ Education captures age’s effect on education. This DAG structure necessitates adjustment for both confounders to identify the causal effects of information sources on fear. We exclude infection counts in the model because we assume that individuals learn about epidemiological conditions exclusively through information sources. In the absence of such sources, individuals would have no access to the actual infection counts. To account for temporal dynamics, we analyze two distinct waves, separated by the date on which infections reached their minimum between the two periods under consideration: May 2021–November 11, 2021, and November 12, 2021–June 2022. Within each wave-specific analysis, we acknowledge that daily observations may exhibit temporal correlation.
Source ATE Source ATE
W1 W2 W1 W2
Scientists 0.24 0.37 Doctors -0.17 0.00
Government 0.18 0.25 Journalists -0.38 -0.25
CDC 0.18 0.33 Religious leaders -1.62 -1.50
Politicians 0.14 0.24 Family and friends -1.90 -1.89
Table 3: ATE of information sources on Fear across two waves using None of the above as the “control” source. W1 and W2 denote the first and second waves of the data.
Specifically, we applied this model—normalizing $\phi_{s^{\prime}}$—to estimate the strength of each direct causal pathway leading to Fear. Our analysis shows that Source is the most influential variable, accounting for 81.8% of the normalized $\phi_{s^{\prime}}$ variance in the first wave and 81.5% in the second—Figures 6.a) and 6.b). This indicates that this variable plays a dominant role in shaping fear responses within the population studied. In comparison, demographic factors such as Age and Education contribute more modestly, accounting for 6.7%-8.8% and 11.5%-9.7% of the variance, respectively, across the two waves. These findings suggest that while individual characteristics do have some impact, the Source variable is the primary driver. Moreover, Figures 6.c) and 6.d) show the influence of each Source, with Family and friends and Religious leaders exerting the greatest influence, while all others have smaller effects—except for Doctors in the first wave, which account for 6.3%.
Finally, we estimate the ATE of different information sources on Fear, using None of the above as the “control” source. Table 3 presents the results. In both waves, the ranking of sources in decreasing order is largely consistent, with the exception of CDC in the second wave, whose effect is higher and closer to that of Scientists. Interestingly, from the first to the second wave, the ATE of each source increases, indicating that during the Omicron wave, people were more afraid when using a source other than the “control” source. Notably, we observe that people are less afraid when obtaining information from the two outliers, Family and friends and Religious leaders, as well as from Journalists, suggesting lower trust in these sources, whereas they are more afraid when obtaining information from more official sources such as Scientists, Government, CDC, and Politicians.
Geographical distribution of sources: U.S. nationwide. Using the definitions introduced in the Data section, we analyze patterns in the use of each singleton information source, independently of Fear, represented by $\hat{\sigma}_{s^{\prime}}$. As before, we begin by analyzing the nine distributions: the Shapiro–Wilk test indicates that none follow a normal distribution. To assess differences—both within and between groups—we used the Kolmogorov–Smirnov and Mann–Whitney U tests. The observed significant differences reinforce the idea that perceived uncertainty varies meaningfully across demographic segments and source groupings.
Based on this evidence, Figure 7 shows the evolution of the fraction of respondents relying on each information source, alongside the official number of infections across the U.S. It is notable that None of the above consistently shows values near zero with little variation and is therefore excluded from the visualization, indicating that few respondents rely on sources outside those listed. Similar to the fear score, respondents’ reliance on information sources closely follows infection trends, showing strong correlation and greater sensitivity to relative increases in cases, except for Politicians, whose pattern remains largely flat and uncorrelated—Figure 8. Some sources exhibit moderate sensitivity. The coefficient of variation varies between approximately $0.25$ for None of the above to around $0.1$ for Doctors and CDC, and is lower for other sources, indicating moderate to low variability.
Figure 7: Temporal evolution of $\hat{\sigma}_{s^{\prime}}$ for each of the nine singleton information sources—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
Respondents’ use of information sources displays strong mutual correlations, as shown in Figure 8, which presents the Spearman correlation coefficients. Within this context, three sources emerge as outliers: Religious leaders—as previously noted—, None of the above, and Politicians. Specifically, Religious leaders and None of the above display negative correlations with all other time series and a positive correlation with each other. This pattern suggests that their trends move in the opposite direction relative to the other sources, indicating that as infection rises, individuals tend to shift away from these less conventional or less trusted sources toward more reliable information channels—an inference supported by their negative correlation with official infection numbers. In contrast, Politicians exhibit negative correlations with Religious leaders, None of the above, and Doctors, while showing a small positive correlation with the other sources. As illustrated in Figure 7, the time series for Politicians remains relatively flat, implying that public reliance on political figures as information providers remains largely stable over time, despite fluctuations in infection severity.
Figure 8: Spearman correlation coefficients among the time series shown in Figure 7.
Finally, Figure 9 displays boxplots illustrating the distribution of the fraction of individuals relying on each singleton information source, stratified by Age and Education groups, alongside the overall distribution aggregated without stratification. In this case as well, we excluded None of the above from the visualization, as it stands out as a clear outlier in both instances. Three clear patterns emerge when moving from $Age_{1}$ to $Age_{3}$ and from $Edu_{1}$ to $Edu_{3}$. First, certain sources maintain a relatively stable or flat level of reliance across these groups, suggesting that demographic factors have little effect on their usage—like Family and friends, Politicians, and None of the above. Second, Religious leaders exhibit a decreasing trend, with younger or less educated groups relying on them more than older or more educated ones. Third, the remaining sources—institutional or expert ones—show increased reliance as Age and Education rise—meaning older or more educated individuals tend to trust these sources more.
Figure 9: Boxplots showing the distribution of the fraction of respondents relying on each information source, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
Geographical distribution of sources: U.S. states. To further understand patterns in the use of information sources, we conduct a geographical analysis across U.S. states. The United States presents a diverse socio-economic landscape, with substantial variation in income levels, educational attainment, and cultural norms across regions. These differences are likely to influence how individuals seek and trust information, particularly in the context of a public health crisis such as the COVID-19 pandemic. By examining the spatial distribution of preferred information sources, we aim to uncover regional trends and disparities that may reflect underlying socio-demographic and economic characteristics. Hence, for each U.S. state, we compute the average reliance on each singleton information source—from $\hat{\sigma}_{s^{\prime}}$—, without stratification by Age or Education, resulting in nine values per state. These averages are then used to cluster the states using K-means, spectral, and hierarchical clustering methods.
Figure 10: Clusters derived from K-means analysis based on average reliance on each information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, without stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections. Misclassified states are overlaid with a contrasting grid. States with a purple border indicate swing states between the 2020 and 2024 U.S. political elections.
Based on the Silhouette score, we determined that two clusters provide the optimal partitioning. Figure 10 illustrates the resulting clusters, with each state colored according to its assigned cluster. Interestingly, by chance, we obtained two clusters that correspond closely to the two major U.S. political affiliations: Democratic and Republican. Specifically, Figure 10.a) compares the derived clusters to the 2020 U.S. political elections results, while Figure 10.b) shows the comparison with the 2024 elections outcomes, both obtained using K-means clustering—similar results were obtained with hierarchical and spectral clustering. States that were misclassified—i.e., assigned to the cluster contrary to their political affiliation in the given year—are marked with a grid overlay in the contrasting cluster color. We observe that only three states were assigned to the opposite cluster in the 2020 elections, while five states were assigned to the opposite cluster in the 2024 elections. Interestingly, these states—except for Alaska—are swing states between the 2020 and 2024 political elections, which makes them inherently more difficult to classify. These results suggest that the clustering based on information source reliance closely reflects established political divisions, with misclassifications primarily occurring in politically competitive swing states. This underscores the sensitivity of information consumption patterns to shifting political landscapes.
Conclusion
Building on behavioral and demographic data, this study explores the factors driving fear of COVID-19 infection, focusing on how information sources and demographics shape emotional responses. We examine how channels such as social media, traditional journalism, government briefings, and interpersonal communication relate to differences in perceived fear across the U.S., analyzing data from May 2021 to June 2022, covering the Delta and Omicron waves.
Recognizing potential biases in digital platforms, particularly toward users with consistent internet access, we contextualize our findings within the socio-economic and geographic diversity of the U.S. Regional disparities in media habits, internet penetration, and education underscore the importance of a geographically nuanced approach to understanding information dynamics and public emotion. Moreover, this study analyzes data exclusively from the U.S. Delphi CTIS survey, which is restricted to American respondents. While this enables detailed causal inference within the U.S. context, the work could be extended to other countries if comparable longitudinal panel data on COVID-19 fear, information sources, and related behaviors become available in the future.
Our analysis shows that COVID-19–related fear associated with the considered information sources is strongly correlated across sources and with the number of confirmed infections. Using causal inference techniques, we quantified the effects of age, education, and information sources on perceived fear. Information sources emerged as the most significant driver, while demographics also shape emotional reactions and preferences for specific types of information, revealing distinct behavioral patterns across age and education groups. These results underscore the pivotal role of the information environment in shaping public fear during health crises. The credibility, framing, and volume of content can amplify or mitigate fear, with misinformation posing a serious threat to risk perception and trust. Ensuring access to verified, reliable sources is therefore critical for informed decision-making and emotional resilience.
Moreover, our findings reveal a strong connection between information ecosystems and political behavior. Clustering analysis showed that patterns of reliance on specific information sources across U.S. states closely aligned with political orientations during the 2020 and 2024 election cycles. These results underscore the broader societal impact of informational environments, demonstrating their influence not only on emotional responses but also on civic attitudes and decision-making. In this context, it becomes imperative that official information channels maintain the highest standards of accuracy and credibility. The dissemination of misinformation—whether intentional or due to negligence—by institutional sources not only distorts public understanding but also undermines democratic processes by shaping political beliefs on falsehoods. Therefore, safeguarding trusted information sources from distortion is crucial for protecting both public health and the integrity of democratic society.
Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes, see the Data and Ethical declaration sections.
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes, the Abstract and the Introduction section state the main contributions of the paper, and we explicitly summarize them in a dedicated Contributions subsection within the Introduction section.
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes, in the Methods section.
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? Yes, in the Data and Conclusion sections.
    5. (e)
      Did you describe the limitations of your work? Yes, in the Data, Methods, an Results sections.
    6. (f)
      Did you discuss any potential negative societal impacts of your work? NA.
    7. (g)
      Did you discuss any potential misuse of your work? NA.
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? NA.
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes.
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? NA
    2. (b)
      Have you provided justifications for all theoretical results? NA
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? NA
    6. (f)
      Have you related your theoretical results to the existing literature in social science? NA
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? Yes, in the Conclusion section.
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? Yes, see the Availability of materials and data section.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? NA
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? NA
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? NA
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? NA
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets, without compromising anonymity
    1. (a)
      If your work uses existing assets, did you cite the creators? Yes, we have cited relevant papers.
    2. (b)
      Did you mention the license of the assets? NA
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? NA
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? Yes, see the Ethical declaration section.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? Yes, the data we used does not contain personally identifiable information or offensive content.
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR? NA
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset? NA
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects, without compromising anonymity
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? NA
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? NA
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? NA
Appendix A: Extended analysis on clustering
In addition to the analysis presented in the Results section—specifically in the Geographical distribution of sources: U.S. states subsection—, we provide additional details on clustering using $\hat{\sigma}_{s^{\prime}}$. We also repeat the analysis with alternative inputs, including $\phi_{s^{\prime}}$ alone and in combination with $\hat{\sigma}_{s^{\prime}}$. In each case, we present the clusters obtained using K-means, noting that similar results were obtained with hierarchical and spectral clustering methods.
Clustering using $\hat{\sigma}_{s^{\prime}}$. Alongside the results shown in the Results section, we incorporate demographic information by computing, for each U.S. state, the average reliance on each singleton information source, grouped by Age, resulting in 27 values per state. As shown in Figure 11, this adjustment reduced misclassifications in the 2024 elections, with only Wisconsin and Alaska assigned to the opposite cluster—but it led to a slight increase in misclassifications for the 2020 elections. A similar improvement is observed when combining information source preferences with Education levels.
Figure 11: Clusters derived from K-means analysis based on average reliance on each singleton information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
The improved classification performance observed for the 2024 elections when incorporating age and education suggests that these demographic variables became more closely associated with political orientation, likely due to shifts in how different age and education groups access and respond to information. In contrast, the reduced accuracy for the 2020 elections indicates that these factors were less informative during that period, possibly due to more uniform voting patterns across demographics or the influence of external events such as the COVID-19 pandemic.
Clustering using $\phi_{s^{\prime}}$. For each U.S. state, we compute the fear score for each singleton information source, $\phi_{s^{\prime}}$, without stratifying by Age or Education, resulting in nine values per state. Figure 12 presents the resulting clusters and compares them with the outcomes of the 2020 and 2024 political elections.
Compared to the clusters obtained using $\hat{\sigma}_{s^{\prime}}$, a larger number of states are misclassified in both elections—specifically, Colorado, Connecticut, Delaware, Illinois, Maine, Massachusetts, Minnesota, New Hampshire, Oregon, Rhode Island, Texas, Vermont, Virginia, and Washington. Notably, swing states are misclassified only in the 2020 elections, but appear in the correct cluster for 2024. Additionally, Alaska is now correctly classified in both elections.
Next, we stratify the data by Age and Education, yielding 27 values per state. Stratifying by Age yields similar results, with the same set of states appearing misclassified. When stratifying by Education, two additional states—New York and Maryland—are misclassified in both the 2020 and 2024 elections. This shift suggests that, for these states, educational differences may influence the relationship between fear scores and political alignment in a manner that diverges from the general pattern observed without stratification by education.
Figure 12: Clusters derived from K-means analysis based on the average respondents’ fear score for each singleton information source, $\phi_{s^{\prime}}$, across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
More broadly, the presence of traditionally Democratic-leaning states among those misclassified indicates that ideological or cultural factors—such as attitudes toward information sources, regional media ecosystems, or local political discourse—might play a role in shaping fear perception patterns that do not align neatly with national political divisions.
Clustering using $\hat{\sigma}_{s^{\prime}}$ and $\phi_{s^{\prime}}$.
Figure 13: Clusters derived from K-means analysis based on the average reliance $\hat{\sigma}_{s^{\prime}}$ on each singleton information source and average respondents’ fear score $\phi_{s^{\prime}}$ for them across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
For each U.S. state, we compute the average reliance $\hat{\sigma}_{s^{\prime}}$ and fear scores $\phi_{s^{\prime}}$ for each singleton information source without stratifying by Age or Education, resulting in 18 values per state. Figure 13 presents the resulting clusters obtained and compares them to the outcomes of the 2020 and 2024 U.S. political elections.
When combining both $\hat{\sigma}_{s^{\prime}}$ and $\phi_{s^{\prime}}$, we observe that five states—Alaska, Georgia, Michigan, Pennsylvania, and Wisconsin—are misclassified for the 2020 elections, while only three states—Alaska, Arizona, and Nevada—are misclassified for the 2024 elections.
Next, we stratify the data separately by Age and Education for both $\hat{\sigma}_{{s^{\prime}}}$ and $\phi_{{s^{\prime}}}$, yielding 54 values per state. Stratifying by Education yields similar results, with the same set of states appearing misclassified. In contrast, when stratifying by Age, only Alaska remains in the wrong cluster for the 2024 elections. This suggests that age-related differences in information consumption and fear perception play a key role in accurately classifying Arizona and Nevada. In these states, incorporating age stratification captures important nuances in how fear and reliance on information sources vary across generations, ultimately aligning their political clustering more accurately.
Interestingly, Wisconsin exhibits the opposite pattern in the 2024 elections. While clustering with $\hat{\sigma}_{s^{\prime}}$ and stratifying by age alone is insufficient for correct classification, incorporating fear scores—or a combination of fear and reliance for each information source—is essential for accurate placement. This suggests that, in Wisconsin, the emotional response to information sources plays a more decisive role in shaping political alignment than demographic factors alone.
Appendix A Ethical declaration
Further details will be provided upon acceptance, in accordance with double-blinded submission requirements.
Appendix B Availability of materials and data
All datasets used in this work will be publicly available. Further details will be provided upon acceptance, in accordance with double-blinded submission requirements.
Appendix C Declaration of competing interest
The authors declare no competing interests.
Appendix D Funding
Further details will be provided upon acceptance, in accordance with double-blinded submission requirements.